a resistance of 60 ohm is immersed in 42kg of water.a current of 7A is passed through it.the rise in temperature of water per minute is

Grade:8

1 Answers

Umakant biswal
7 years ago
@ m gayathri 
i am giving you the proper approach to solve this question . 
heat transfered to water is i^2Rt = 7^2 * 60 * 60
this is equal to m s delta theta. 
m is mass of water
specific heat of water is 4.186 joule/gram °C. 
find delta theta. 
by equating both. 
i^2 R t = ms delta theta, m is mass, s is specific heat, t is time. 
I THING NOW U CAN SOLVE THE QUESTION 
IF NOT THEN LET ME KNOW . 
all the best ..